OpenStudy (anonymous):

x^2+ 4xy + 4y^2 +6xz+ 12yz +9z^2 CAN YOU HELP FIND THE SQUARE ROOT? I HAVE NO IDEA HOW TO GET STARTED!!!

OpenStudy (anonymous):

1st find like terms

OpenStudy (anonymous):

wait whats the question

OpenStudy (anonymous):

\[\sqrt{x^2 + 4xy +4y^2 +6xz+12yz+9z^2 }\]

OpenStudy (loser66):

=x + 3z +2y

OpenStudy (loser66):

I walk you through this stuff:

OpenStudy (anonymous):

ok that would be great!

OpenStudy (loser66):

is it unclear to you for the last step?

OpenStudy (anonymous):

a little bit on the last step.

OpenStudy (loser66):

\[\sqrt{((x+3z)+(2y))^2}= (x+3z)+(2y)= x +3z +2y\]

OpenStudy (loser66):